# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ Appendix B Errors and convergence in experimental current measurements
It is particularly pertinent to check for convergence in Experiments 2 and 3, where waves are present during the measurements, with instantaneous orbital velocities which cause differences from frame to frame which far exceed $\Delta U$. In Experiment 2 individual snapshots of the velocity field were taken at low frequency. In figure 8(c) we show the absolute difference (L1 norm) between the average over the first $n$ snapshots, and the full set of $2000$ snapshots. Although convergence is slow as can be expected, averages vary by less than $1$ mm/s after about $n=1200$ frames, and it is clear that a longer time series would not significantly alter the results. Likewise, convergence with increasing number of repeated measurements in Experiment 3 is shown for case 3.B, representative also of case 3.A. The average no longer fluctuates significantly after about $20$ out of the $32$ $50$-second intervals.
@@ Appendix C Separating waves from turbulence
Figure 9: Wave-turbulence decomposition of the streamwise velocity field $\check{u}$ using POD. (a) normalized mode energy $\lambda_{n}$ for a single ensemble of case 3.A.2. (b) vertical slice of mode 1 beneath a wave peak ($x=x_{\mathrm{p}}$). A factor $C$ is used for normalisation. (c) Fourier transform of the temporal coefficients of mode 1 and the wave probe signal. (d) Spectrogram of the streamwise velocity at each depth coordinate; see equation (7). (e) Snapshots of the decomposed signal. From left to right in panels (d) and (e): $\check{u}(\mathbf{x},t)$, $\tilde{u}(\mathbf{x},t)$, and $u(\mathbf{x},t)^{\prime}$
For the analysis in figure 6 a triple decomposition according to (7) is needed. The task is nontrivial as demonstrated by many authors in the past (see the thorough review by Chávez-Dorado et al., 2025), and particularly challenging for our cases 3.A and 3.B where phase information is not available. Standard methods such as phase-conditioned averaging (PhCA), and other methods such as Dynamic Mode Decomposition (Chávez-Dorado et al., 2025), Synchrosqueeze Wavelet Transform (Perez et al., 2020) and Empirical Mode Decomposition (Peruzzi et al., 2021) cannot be employed. In contrast, measurements from Experiment 3 have a spatially similar plane of measurement and field of view, and are resolved in time so that both PhCA and POD can be used, allowing us to validate POD for cases 3.A and 3.B.
@@ C.1 Proper Orthogonal Decomposition
Here we briefly present the principle of POD while referring to specialised literature for details (e.g., Berkooz et al., 1993; Taira et al., 2017). In short, after subtracting the mean velocity, the remaining field $\check{u}(\mathbf{x},t)$ is decomposed into $N$ orthogonal spatial modes, $\phi_{n}(\mathbf{x})$, and their respective temporal coefficients, $a_{n}(t)$, as
$\check{u}(\mathbf{x},t)=\sum_{n=1}^{N}a_{n}(t)\phi_{n}(\mathbf{x})$ (37)
where $N$ is the number of measured PIV fields, and each mode has an associated ‘energy’ $\lambda_{n}$. POD is calculated using the method of snapshots.
Figure 9 illustrates POD performed on case 3.A. Panel (a) shows the POD mode energy distribution for a single ensemble. Note how the two highest-energy modes, $\lambda_{1}$ and $\lambda_{2}$ contribute about $90\%$ of the energy in the flow indicating that POD identifies a low-rank structure (see, e.g., Taira et al., 2017). It transpires that according to criteria we will return to, these two modes can be identified as containing the wave motion $\tilde{{\boldsymbol{u}}}$.
@@
A more quantitative test of the method’s performance is Figure 9 (d) which shows the power spectral density as a function of depth, indicating where in the water column, and for which frequencies, kinetic energy has been extracted. The peak in energy around the wave frequency is nearly exclusively present in the spectrum of $\tilde{u}(z)$ (central panel). A small amount of energy is present also at frequencies far below $\omega_{0}$, which is expected due to minor fluctuations in the mean velocity, subharmonic waves, and low-frequency sloshing modes in the tank. Finally, for illustrative purposes, a decomposed signal from a single PIV frame is displayed in Figure 9(e).
@@ C.2 Phase-conditioned average
Figure 10: Wave-turbulence decomposition using PhCA from measurements of case 3.A.2. (a) Sample of the wave phase field, $\Phi(x,t)$, showing the phase variation across the spatial domain just beneath the wave trough. (b) Phase-resolved average velocity, $\hat{u}(\Phi,z)$. (c) and (d): Same as figure 9(d) and (e), respectively, for PhCA instead of POD.
Variations of the PhCA method have been employed in wave/turbulence research for a long time. Following the procedure of (Buckley & Veron, 2017), the wave motion is obtained by extracting the mean velocity field for each value of the depth and the phase of the wave Buckley & Veron (2017), denoted $\hat{u}(\Phi,z)$.
The wave phase, $\Phi(x,t)$, is extracted from the analytic signal of the velocity at a depth layer just beneath the wave trough. At each position, the analytic signal is acquired from the Hilbert transform in the temporal direction (Melville, 1983). A sample wave phase field is depicted in 10(a). The phase is divided into bins, and the phase-conditioned average, $\hat{u}(\Phi,z)$ is calculated by taking the average of the mean-subtracted velocity $\check{u}(\mathbf{x},t)$ in each phase bin. A sample of the resulting phase-resolved average, $\hat{u}(\Phi,z)$, is shown in Figure 10(b). The wave component $\hat{u}(\mathbf{x},t)$ is reconstructed from $\Phi(x,t)$ and $\hat{u}(\Phi,z)$. Figure 10(c) shows the power spectral density as a function of depth for comparison with figure 9(d), and figure 10(d) shows the same snapshot of the velocity field as figure 10(e).
Figure 11: Difference between triple-decomposed streamwise velocity fields for two different snapshots with approximately the same phase from case G.2, streamwise velocity component (waves moving left to right, current moving right to left). Left panels: full mean-subtracted velocity field $\check{u}$; middle panels: difference in wave velocities; right panels: difference in turbulent velocities. Top row: error due to PhCA amplitude error; bottom row: error due to PhCA phase error.
@@
We compare first the qualitative information in the shapshots of the decomposed velocity fields. In the case of POD in figure 9(e), some tiny velocity fluctuations remain in the supposed wave component in the former, faintly visible in the lower half of the central panel. These fluctuations are not correlated to the turbulent flow field and are an inescapable artifact of the mode decomposition procedure. PhCA by construction produces a field $\tilde{{\boldsymbol{u}}}$ which is smooth, as seen in figure 10(d). The snapshots of the extracted $u^{\prime}$ in the rightmost panels are hardly possible to distinguish by eye, but the quantitative comparison shows that more energy is ascribed to turbulence (and less to waves) for PhCA than POD.
An even closer comparison reveals that there are three main types of discrepancies between the two methods. The wave velocity $\tilde{u}$ also contains considerable signal at frequencies well below $\omega_{0}$; surface waves of such low frequencies are considerably longer than the field of view and PhCA cannot detect them.
Second, we regard the depth-spectrograms in figures 9(d) and 10(c), focussing on frequencies near the carrier frequency where the spectra have a very pronounced peak. The “wave residue” in the spectrum near this frequency after waves have been removed is significantly smaller for POD than for PhCA, indicating that the latter leaves a larger ‘wave residue’ in the identified turbulence field ${\boldsymbol{u}}^{\prime}$. PhCA produces a more narrowband $\tilde{{\boldsymbol{u}}}$ with hardly any signal outside of this peak, whereas the POD wave velocities produce a broader wave spectrum. A look at the difference between the POD and PhCA wave fields, as in the middle panels in figure 11, reveals that the wave-signal assigned to turbulence is due to slight inaccuracies of PhCA to determine phase and amplitude. The top row (Frame 81) shows an amplitude mismatch (the difference field is in phase with the wave), the bottom row (Frame 201) shows a phase mismatch (difference field approximately $\pi/2$ out of phase with the wave).
@@ Appendix D Reynolds stress measurements
For deeper consideration of our results such as figure 6, it is instructive to regard the Reynolds stresses measured for the cases 2.A3.B with regular waves. These are shown in figure 12; panels (a)–(f) show the stresses themselves while the ratio $\overline{u^{\prime}u^{\prime}}/\overline{w^{\prime}w^{\prime}}$ which enters in (13) is shown in panels (g)–(l). The vertical behaviours of $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ are qualitatively similar to those found numerically by Fujiwara & Yoshikawa (2020) (their figure 8 — we cannot reliably capture the region very close to the surface where $\overline{w^{\prime}w^{\prime}}$ must by necessity tend to zero.)

### v2

@@ Appendix B Errors and convergence in experimental current measurements
It is particularly pertinent to check for convergence in Experiments 2 and 3, where waves are present during the measurements, with instantaneous orbital velocities which cause differences from frame to frame which far exceed $\Delta U$. In Experiment 2 individual snapshots of the velocity field were taken at low frequency. In figure 8(c) we show the absolute difference (L1 norm) between the average over the first $n$ snapshots, and the full set of $2000$ snapshots. Although convergence is slow as can be expected, averages vary by less than $1$ mm/s after approximately $n=1200$ frames, and it is clear that a longer time series would not significantly alter the results. Likewise, convergence with increasing number of repeated measurements in Experiment 3 is shown for case 3.B, representative also of case 3.A. The average no longer fluctuates significantly after approximately $20$ out of the $32$ $25$-second intervals.
@@ Appendix C Separating waves from turbulence
Figure 9: wave–turbulence decomposition of the streamwise mean-subtracted velocity field $u_{\mathrm{wt}}$ using POD. (a) Normalised mode energy $\lambda_{n}$ for a single ensemble of case 3.A.2. (b) Vertical slice of mode 1 beneath a wave peak ($x=x_{\mathrm{p}}$). A factor $C$ is used for normalisation. (c) Power spectral density (PSD) of the temporal coefficients of mode 1 and the wave probe signal. (d) Spectrogram of the streamwise velocity at each depth coordinate; see equation (7). (e) Snapshots of the decomposed signal. From left to right in panels (d) and (e): $_{\mathrm{wt}}, u~\tilde{u}$, and $u^{\prime}$. A supplementary movie illustrating wave–current decomposition of our data is available at https://doi.org/10.1017/jfm.2026.11163.
For the analysis in figure 6 a triple decomposition according to (7) is needed. The task is nontrivial as demonstrated by many authors in the past (see the thorough review by Chávez-Dorado et al., 2025), and particularly challenging for our cases 3.A and 3.B where phase information is not available. Standard methods such as phase-conditioned averaging (PhCA), and other methods such as Dynamic Mode Decomposition (Chávez-Dorado et al., 2025), Synchrosqueeze Wavelet Transform (Perez et al., 2020) and Empirical Mode Decomposition (Peruzzi et al., 2021) cannot be employed for data from Experiment 2. In contrast, measurements from Experiment 3 have a spatially similar plane of measurement and field of view, and are resolved in time so that both PhCA and POD can be used, allowing us to validate POD for cases 3.A and 3.B.
@@ C.1 Proper Orthogonal Decomposition
Here we briefly present the principle of POD while referring to specialised literature for details (e.g., Berkooz et al., 1993; Taira et al., 2017). In short, after subtracting the mean velocity $\bar{u}(z)$, each component of the remaining field which we denote ${\boldsymbol{u}}_{\mathrm{wt}}={\boldsymbol{u}}-\bar{{\boldsymbol{u}}}=\tilde{{\boldsymbol{u}}}+{\boldsymbol{u}}^{\prime}$ is decomposed into $N$ orthogonal spatial modes, $\phi_{n}(\mathbf{x})$, and their respective temporal coefficients, $a_{n}(t)$, as
$u_{\mathrm{wt}}(\mathbf{x},t)=\sum_{n=1}^{N}a_{n}(t)\phi_{n}(\mathbf{x})$ (40)
($u$ is understood to be either velocity component) where $N$ is the number of measured PIV fields, and each mode $n$ has an associated ‘energy’ $\lambda_{n}$. We calculate the POD using the method of snapshots.
Figure 9 illustrates POD performed on case 3.A. Panel (a) shows the POD mode energy distribution for a single ensemble. Note how the two highest-energy modes, $\lambda_{1}$ and $\lambda_{2}$ contribute approximately $90\%$ of the energy in the flow indicating that POD identifies a low-rank structure (see, e.g., Taira et al., 2017). It transpires that according to criteria we will return to, these two modes can be identified as containing the wave motion $\tilde{{\boldsymbol{u}}}$.
@@
A more quantitative test of the method’s performance is Figure 9 (d) which shows a depth-wise power spectral density. By calculating the Fourier transform in time for points at the same $z$, the plot visualises where in the water column, and for which frequencies, kinetic energy has been extracted. The peak in energy around the wave frequency is nearly exclusively present in the spectrum of $\tilde{u}(z)$ (central panel). A small amount of energy is present also at frequencies far below $\omega_{0}$, which is expected due to minor fluctuations in the mean velocity, subharmonic waves, and low-frequency sloshing modes in the tank. Finally, for illustrative purposes, a decomposed signal from a single PIV frame is displayed in Figure 9(e).
@@ C.2 Phase-conditioned average
Figure 10: wave–turbulence decomposition of the streamwise velocity component using PhCA from measurements of case 3.A.2. (a) Sample of the wave phase field, $\Phi(x,t)$, showing the phase variation across the spatial domain just beneath the wave trough. (b) Phase-resolved average velocity, $\tilde{u}_{\mathrm{ph}}(\Phi,z)$. Panels (c) and (d) show the same as figures 9(d) and (e), respectively, for PhCA instead of POD.
Variations of the PhCA method have been employed in wave/turbulence research for a long time. Following the procedure of (Buckley & Veron, 2017), the wave motion is obtained from $u_{\mathrm{wt}}$ by extracting the mean velocity field for each value of the depth and the phase of the wave, denoted $\tilde{u}_{\mathrm{ph}}(\Phi,z)$.
The wave phase, $\Phi(x,t)$, is extracted from the analytic signal of the velocity at a depth layer just beneath the wave trough. At each position, the analytic signal is acquired from the Hilbert transform in the temporal direction (Melville, 1983). A sample wave phase field is depicted in 10(a). The phase is divided into bins, and the phase-conditioned average, $\tilde{u}_{\mathrm{ph}}(\Phi,z)$ is calculated by taking the average of the mean-subtracted velocity $u_{\mathrm{wt}}({\boldsymbol{x}},t)$ in each phase bin. A sample of the resulting phase-resolved average, $\tilde{u}_{\mathrm{ph}}(\Phi,z)$, is shown in Figure 10(b). The wave component of the velocity field is taken to be $\tilde{u}(\mathbf{x},t)=\tilde{u}_{\mathrm{ph}}(\Phi(x,t),z)$. Figure 10(c) shows the power spectral density as a function of depth for comparison with figure 9(d), and figure 10(d) shows the same snapshot of the velocity field as figure 10(e).
Figure 11: Difference between triple-decomposed streamwise velocity fields for two different snapshots with approximately the same phase from case 3.A.2, streamwise velocity component (waves moving left to right, current moving right to left). Left panels: full mean-subtracted velocity field $u_{\mathrm{wt}}$; middle panels: difference in wave velocities; right panels: difference in turbulent velocities. Top row: error due to PhCA amplitude error; bottom row: error due to PhCA phase error.
@@
We compare first the qualitative information in the snapshots of the decomposed velocity fields. In the case of POD in figure 9(e), some tiny velocity fluctuations remain in the supposed wave component, faintly visible in the lower half of the central panel. These fluctuations are not correlated to the turbulent flow field and are an inescapable artefact of the mode decomposition procedure. PhCA by construction produces a field $\tilde{{\boldsymbol{u}}}$ which is smooth, as seen in figure 10(d). The snapshots of the extracted $u^{\prime}$ in the rightmost panels of figures 9(e) and 10(d) are hardly possible to distinguish by eye, but the quantitative comparison shows that more energy is ascribed to turbulence (and less to waves) for PhCA than POD.
An even closer comparison reveals that there are two main types of discrepancies between the two methods. First, the wave velocity $\tilde{u}$ also contains considerable signal at frequencies well below $\omega_{0}$; surface waves of such low frequencies are considerably longer than the field of view and PhCA cannot detect them. The second difference is clearly seen when regarding in the depthwise spectra of $u^{\prime}$ in figures 9(d) and 10(c), when we consider the frequencies near the carrier-wave frequency $f_{0}$ (equal to $1.78$ Hz in this example) where the spectrum of the whole field $u_{\mathrm{wt}}$ has a very pronounced peak. Both methods leave some residue of the wave in the turbulent signal after waves have been removed, but it is significantly smaller for POD than for PhCA, indicating that the latter leaves a larger ‘wave residue’ in the identified turbulence field $\boldsymbol{u}$$^{\prime}$. PhCA produces a more narrowband $\tilde{{\boldsymbol{u}}}$ with hardly any signal outside of this peak, whereas the POD wave velocities produce a broader wave spectrum.
A look at the difference between the POD and PhCA wave fields, as in the middle panels in figure 11, reveals that the wave-signal assigned to turbulence is due to slight inaccuracies of PhCA to determine phase and amplitude. The top row (Frame 81) shows an amplitude mismatch (the difference field is in phase with the wave), and the bottom row (Frame 201) shows a phase mismatch (difference field approximately $\pi/2$ out of phase with the wave).
@@ Appendix D Reynolds stress measurements
For deeper consideration of our results such as figure 6, it is instructive to regard the Reynolds stresses measured for the cases 2.A3.B with regular waves. These are shown in figure 12; panels (a)–(f) show the stresses themselves while the ratio $\overline{u^{\prime}u^{\prime}}/\overline{w^{\prime}w^{\prime}}$ which enters in (13) is shown in panels (g)–(l). The vertical behaviours of $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ are qualitatively similar to those found numerically by Fujiwara & Yoshikawa (2020) (their figure 8; we cannot reliably capture the region very close to the surface where $\overline{w^{\prime}w^{\prime}}$ must by necessity tend to zero.)